• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 03:58:53 on April18,2007

    Entry number 199452

    keyword=CODA problems

    during run 2609, we lost about 40 minutes because CODA stopped reading events, even though the norm. scalers were still fine, beam was still on, CODA was still running, and communication with all the ROCS and TS was still good, so we didn't notice anything was happening until we noticed that it had been a long time since a run had ended. So we tried to end the run, which failed. Therefore, we did a "startcoda" on adaql2 and everything reset just fine. We started run 2610 and so far are running without problems.